# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v2

@@ 2.4 Main Result
The main contribution of our work is to rigorously show that the definitions for linear stability/instability, heuristically derived in section 2.3, characterize the supports of $X^{\operatorname{GD/SGD}}_{\lim}$ and thus characterize the qualitative implicit bias of gradient descent/stochastic gradient descent. In our derivation we made some assumptions, which can be formalized as follows.
@@
is linearly independent. In particular, $\mathcal{M}$ is an embedded sub-manifold of $\mathbb{R}^{D}$.
@@
The initial distribution $\nu$ is equivalent to Lebesgue measure, i.e.
@@
Recall that, by an argument by Cooper (2021), the Hypothesis (H1) is fulfilled for generic training data, while (H2) holds for the most common choice of initial distribution.
Furthermore, the heuristic argument presented in the previous section relied on the assumption that with probability one the optimization algorithms do not reach $\mathcal{M}$ in a finite number of steps. Hypothesis (H2) ensures that, with probability one, the algorithms are not initialized on $\mathcal{M}$. However, the setting presented so far is too general to exclude that the optimization algorithms reach $\mathcal{M}$ after any positive finite number of iterations with positive probability. Thus we require an additional assumption. From now on, consider a fixed learning rate $\eta>0$.
@@
In particular, a continuously differentiable map $\mathcal{G}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is non-singular, if its Jacobian is invertible Lebesgue almost everywhere. Whether this is true for the maps $\mathcal{G}_{\eta},\mathcal{G}_{\eta,1},\dots,\mathcal{G}_{\eta,N}$ depends on the network function $\mathfrak{F}$. Yet it is reasonable to assume that this should be satisfied for the common neural network architechtures, at least for almost every learning rate $\eta$.
@@
Suppose (H1), (H2) and (H3) are satisfied. Let $x^{*}\in\mathcal{M}$.
@@
The analogous result for stochastic gradient descent requires some extra assumptions on the global minimum $x^{*}$ in question.
@@
A global minimum $x^{*}\in\mathcal{M}$ is said to be regular, if
@@
Note that almost every family of vectors $(\nabla_{x}\mathfrak{F}(x^{*},y_{i}))_{i\in[N]}$ will satisfy these conditions. Thus it is reasonable to assume that, for most network functions $\mathfrak{F}$, almost every global minimum $x^{*}\in\mathcal{M}$ is regular. With this definition in hand, we get the following result for SGD.
@@
Suppose (H1), (H2) and (H3) are satisfied. Let $x^{*}\in\mathcal{M}$ be regular.
@@
Both Theorem A and Theorem B fail to address the case $\mu(x^{*})=0$, respectively $\lambda(x^{*})=0$. Since the support of $X_{\lim}^{GD/SGD}$ is closed by definition and $\mu:\mathcal{M}\to\mathbb{R}$ is continuous, most points with $\mu(x^{*})=0$ should have global minima $x^{\prime}\in\mathcal{M}$ with $\mu(x^{\prime})<0$ nearby and thus be in the support of $X_{\lim}^{GD}$. Making this argument rigorous would require more precise knowledge of the network function $\mathfrak{F}$ and is outside the scope of this paper. Continuity of $\lambda$ is a more subtle issue. In general, $\lambda:\mathcal{M}\to\mathbb{R}$ is only upper semi-continuous [^9] This can readily be seen from the fact that it is defined as an infimum of continuous function.. However, a recent result by Avila et al. (2023) shows that $\lambda$ is continuous at all points $x^{*}$, where the matrices $\mathcal{G}^{\prime}_{\eta,1}(x^{*}),\dots,\mathcal{G}^{\prime}_{\eta,N}(x^{*})$ are all invertible. This is, in particular, the case for all regular $x^{*}\in\mathcal{M}$.
@@ 2.5 Relation to the Neural Tangent Kernel
where $G_{x^{*},[i]}$ denotes the matrix $G_{x^{*}}$ with every but the $i$-th row set to zero. It should be noted that $S_{x^{*}}$ is not a square matrix and thus not invertible as a matrix. However as an isomorphism from $\mathbb{R}^{N}$ to $\mathcal{N}(x^{*})$ it can be inverted and we let $S_{x^{*}}^{-1}\in\mathbb{R}^{N\times D}$ denote the matrix associated to this inverse isomorphism. It satisfies $S^{-1}_{x^{*}}S_{x^{*}}=\mathds{1}_{N}$ and $(S_{x^{*}}S^{-1}_{x^{*}})|_{\mathcal{N}(x^{*})}=\mathds{1}_{D}|_{\mathcal{N}(x^{*})}$. Thus the restrictions of the Jacobians of $\mathcal{G}_{\eta}$ and $\mathcal{G}_{\eta,i}$ to $\mathcal{N}(x^{*})$ can be expressed as
@@
From this we get $\rho_{\operatorname{Spec}}(\mathcal{G}^{\prime}_{\eta}(x^{*})|_{\mathcal{N}(x^{*})})=\rho_{\operatorname{Spec}}(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}})$ and, using the symmetry of $(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}})$, also
$\mu(x^{*})=\rho_{\operatorname{Spec}}\left(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}}\right)=\left\|\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}}\right\|.$ (2.18)
For $\lambda(x^{*})$ an analogous statement can be obtained. Indeed, we have the estimates
@@
Together they imply
@@
This observation is useful for two reasons. On the one hand it connects the central quantities in the present work to an ongoing line of research of neural networks in the infinite width limit (see Jacot et al. (2018)). On the other hand, it turns out that the expressions (2.18) and (2.19) are nicer to work with than the corresponding expressions in the previous section and will be exclusively used in the proofs of the main theorems.
@@ 2.6 Possible Extensions and Outlook
While our main results are general in the sense that they require only very weak assumptions on the network function $\mathfrak{F}$, we want to point out that we only consider scalar regression problems and training with gradient descent or stochastic gradient descent with step size 1. As mentioned above, we have made these restrictions to simplify the already extensive proofs as much as possible. The main results can be extended, with minor modifications, to a more general setting.
In practice SGD is usually implemented with so-called mini-batches to allow for parallel computations. For SGD with mini-batches of size $1\leq B\leq N$, the iterative step (2.8) is replaced by
@@
where $G_{x^{*},[\Xi]}$ denotes the Gram matrix $G_{x^{*}}$ of the neural tangent kernel with all rows whose index is not in $\Xi$ set to 0. In order to derive a version of Theorem B it is necessary to adapt the notion of regular global minima (cf. Definition 2.2) to ensure that the analog of Lemma 3.9 in the proof below still remains valid. We leave this as a problem for future work. The rest of the proof of Theorem B can be extended to mini-batch SGD with only minor changes. Such an extension of Theorem B could help to better understand the impact of the learning rate and the mini-batch size on generalization properties (see e.g. Hoffer et al. (2017); Goyal et al. (2017); Keskar et al. (2017) for numerical studies).
@@
Furthermore it would be interesting to extend our analysis to other versions of stochastic gradient descent such as SGD with momentum (cf. Rumelhart et al. (1986)) or Adam (cf. Kingma and Ba (2014)). The derivations in section 2.3 can be extended to these algorithms by linearizing the iteration steps around fixed points where all moment terms are zero. In the case of Adam, the parameter $\epsilon$ appearing in the numerator of the final update step is usually chosen extremely small (Kingma and Ba (2014) suggest $\epsilon=10^{-8}$). Thus the linearization only approximates the dynamics of the actual algorithm in a vanishingly small neighborhood and it is questionable whether it is still meaningful for the dynamics of the algorithm.
@@
Our analysis can be extended to multi-dimensional regression problems with ground truth function $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}^{\tilde{d}}$ and network function $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}^{\tilde{d}}$. In fact one can interpret such a multi-dimensional regression problem as a scalar regression problem with ground truth function $\hat{f}:\mathbb{R}^{d}\times[\tilde{d}]\to\mathbb{R}$ and network function $\mathfrak{F}:\mathbb{R}^{D}\times(\mathbb{R}^{d}\times[\tilde{d}])\to\mathbb{R}$. This has the effect of splitting one training example $(y,z)\in\mathbb{R}^{d}\times\mathbb{R}^{\tilde{d}}$ into $\tilde{d}$ training examples $((y,1),z_{1}),\dots,((y,\tilde{d}),z_{\tilde{d}})\in(\mathbb{R}^{d}\times[\tilde{d}])\times\mathbb{R}$. A SGD step with the training example $(y,z)$ now corresponds to an SGD step with the mini-batch $((y,1),z_{1}),\dots,((y,\tilde{d}),z_{\tilde{d}})$. Similarly to mini-batch SGD, the main challenge in extending Theorem B to multi-dimensional regression problems is to adapt the notion of a regular global minimum (cf. Definition 2.2) in such a way that Lemma 3.9 still remains valid for multi-dimensional regression.
@@
Unfortunately our results can not be easily extended to classification problems. For a classification task with $k$ classes, it is common to encode the training data as pairs $(y,z)\in\mathbb{R}^{d}\times\mathbb{R}^{k}$, where $z=e_{i}$ is the $i$-th unit vector, where $i\in[k]$ is the class $y$ belongs to. Then one considers a network function $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}^{k}$, which is then post-composed with the softmax function $\operatorname{softmax}_{\beta}:\mathbb{R}^{k}\to[0,1]^{k}$ given by
@@
where $\beta>0$ is some positive real number commonly referred to as inverse temperature. For training, the most common loss function is cross-entropy loss given by
@@
In order to have $\ell(\bar{z},z)=0$, one necessarily needs $\bar{z}=z$. However, the softmax functions outputs are in $(0,1)^{k}$ for any finite input values. Thus it is not possible to reach training error zero and $\mathcal{M}=\emptyset$.
Finally, it should be mentioned that our analysis assumes that the model is trained to convergence. In practice, it has been observed that stopping the algorithm early can improve generalization. While the global minima with $\lambda(x^{*})>0$ are asymptotically unstable, the the finite time Lyapunov exponents
@@
can, even for large $n$, still be negative with small, but positive probability. This effect is captured by a large deviation principle (cf. Arnold and Kliemann (1987)), the rate function of which depends on the moment Lyapunov exponents (see Definition A.2).

### v3

@@ 2.4 Main Result
The main contribution of our work is to rigorously show that the definitions for linear stability/instability, heuristically derived in Section 2.3, characterize the supports of $X^{\operatorname{GD/SGD}}_{\lim}$ and thus characterize the qualitative implicit bias of gradient descent/stochastic gradient descent. In our derivation, we made some assumptions which can be formalized as follows.
@@
is linearly independent. In particular, $\mathcal{M}$ is an embedded submanifold of $\mathbb{R}^{D}$.
@@
The initial distribution $\nu$ is equivalent to the Lebesgue measure, i.e.
@@
Recall that, by an argument by Cooper (2021), Hypothesis (H1) is fulfilled for generic training data, while (H2) holds for the most common choice of initial distribution.
Furthermore, the heuristic argument presented in the previous section relied on the assumption that with probability one the optimization algorithms do not reach $\mathcal{M}$ in a finite number of steps. Hypothesis (H2) ensures that, with probability one, the algorithms are not initialized on $\mathcal{M}$. However, the setting presented so far is too general to exclude that the optimization algorithms reach $\mathcal{M}$ after any positive finite number of iterations with positive probability. Thus, we require an additional assumption. From now on, consider a fixed learning rate $\eta>0$.
@@
In particular, a continuously differentiable map $\mathcal{G}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is non-singular if its Jacobian is invertible Lebesgue almost everywhere. Whether this is true for the maps $\mathcal{G}_{\eta},\mathcal{G}_{\eta,1},\dots,\mathcal{G}_{\eta,N}$ depends on the network function $\mathfrak{F}$. Yet, it is reasonable to assume that this should be satisfied for the common neural network architectures, at least for almost every learning rate $\eta$.
@@
Suppose (H1), (H2), and (H3) are satisfied. Let $x^{*}\in\mathcal{M}$.
@@
The analogous result for stochastic gradient descent requires some additional assumptions on the global minimum $x^{*}$ in question.
@@
A global minimum $x^{*}\in\mathcal{M}$ is said to be regular if
@@
Note that almost every family of vectors $(\nabla_{x}\mathfrak{F}(x^{*},y_{i}))_{i\in[N]}$ will satisfy these conditions. Thus, it is reasonable to assume that, for most network functions $\mathfrak{F}$, almost every global minimum $x^{*}\in\mathcal{M}$ is regular. Together, the conditions (i) and (ii) in Definition 2.2 imply that the semigroup generated by the operators $\mathcal{G}^{\prime}_{\eta,1}(x^{*}),\dots,\mathcal{G}^{\prime}_{\eta,N}(x^{*})$ has nice algebraic properties, namely being contracting and strongly irreducible (cf. Definition 3.8 below). These properties are well established in the literature on random-matrix products (see, e.g. Bougerol and Lacroix, 1985) and allow us to use tools from that theory which are crucial for our proof. With this definition in hand, we get the following result for SGD.
@@
Suppose (H1), (H2), and (H3) are satisfied. Let $x^{*}\in\mathcal{M}$ be regular.
@@
Both Theorem A and Theorem B do not address the case $\mu(x^{*})=0$, respectively $\lambda(x^{*})=0$. Since the support of $X_{\lim}^{GD/SGD}$ is closed by definition and $\mu:\mathcal{M}\to\mathbb{R}$ is continuous, most points with $\mu(x^{*})=0$ should have global minima $x^{\prime}\in\mathcal{M}$ with $\mu(x^{\prime})<0$ near them and therefore be in the support of $X_{\lim}^{GD}$. Making this argument rigorous would require a more precise knowledge of the network function $\mathfrak{F}$ and is beyond the scope of this paper. The continuity of $\lambda$ is a more subtle issue. In general, $\lambda:\mathcal{M}\to\mathbb{R}$ is only upper semi-continuous [^9] This can readily be seen from the fact that it is defined as an infimum of continuous function.. However, a recent result of Avila et al. (2023) shows that $\lambda$ is continuous at all points $x^{*}$, where the matrices $\mathcal{G}^{\prime}_{\eta,1}(x^{*}),\dots,\mathcal{G}^{\prime}_{\eta,N}(x^{*})$ are all invertible. This is, in particular, the case for all regular $x^{*}\in\mathcal{M}$.
@@ 2.5 Relation to the Neural Tangent Kernel
where $G_{x^{*},[i]}$ denotes the matrix $G_{x^{*}}$ with every row but the $i$-th set to zero. It should be noted that $S_{x^{*}}$ is not a square matrix and is thus not invertible as a matrix. However, as an isomorphism from $\mathbb{R}^{N}$ to $\mathcal{N}(x^{*})$ it can be inverted, and we let $S_{x^{*}}^{-1}\in\mathbb{R}^{N\times D}$ denote the matrix associated with this inverse isomorphism. It satisfies $S^{-1}_{x^{*}}S_{x^{*}}=\mathds{1}_{N}$ and $(S_{x^{*}}S^{-1}_{x^{*}})|_{\mathcal{N}(x^{*})}=\mathds{1}_{D}|_{\mathcal{N}(x^{*})}$. Thus, the restrictions of the Jacobians of $\mathcal{G}_{\eta}$ and $\mathcal{G}_{\eta,i}$ to $\mathcal{N}(x^{*})$ can be expressed as
@@
From this we get $\rho_{\operatorname{Spec}}(\mathcal{G}^{\prime}_{\eta}(x^{*})|_{\mathcal{N}(x^{*})})=\rho_{\operatorname{Spec}}(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}})$ and, using the symmetry of $(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}})$, we also get
$\mu(x^{*})=\log\left[\rho_{\operatorname{Spec}}\left(\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}}\right)\right]=\log\left\|\mathds{1}_{N}-\frac{\eta}{N}G_{x^{*}}\right\|.$ (2.18)
An analogous statement can be obtained for $\lambda(x^{*})$. Indeed, we have the estimates
@@
Together, they imply
@@
This observation is useful for two reasons. On the one hand, it connects the central quantities in the present work to an ongoing line of research of neural networks in the infinite-width limit (see Jacot et al., 2018). Additionally, it turns out that the expressions (2.18) and (2.19) are nicer to work with than the corresponding expressions in the previous section and will be used in the proofs of the main theorems. Finally, these expressions are well defined for any $x\in\mathbb{R}^{D}$ and not only for global minima $x^{*}\in\mathcal{M}$. This is especially useful for the potential empirical studies suggested in the following section.
2.6 Edge of Stability and Empirical Studies
While the present work is purely theoretical, the Lyapunov exponent $\lambda(x^{*})$ of a global minimum $x^{*}$ is also an interesting object for empirical studies. As mentioned in the introduction, a highly interesting study by Cohen et al. (2020) observed that training with gradient descent undergoes two distinct phases. At initialization, the Hessian of the loss landscape satisfies $\|\operatorname{Hess}\mathcal{L}(X_{0}^{\operatorname{GD}})\|<\frac{2}{\eta}$. During the early stages of training with gradient descent, the parameters $X_{n}^{\operatorname{GD}}$ move to progressively sharper regions of the loss landscape, that is, the operator norm of the Hessian of the loss function increases. During this phase, called progressive sharpening, the loss $\mathcal{L}(X_{n}^{\operatorname{GD}})$ decreases monotonically. Once $\|\operatorname{Hess}\mathcal{L}(X_{n}^{\operatorname{GD}})\|$ reaches the threshold $\frac{2}{\eta}$, progressive sharpening stops and training continues in the so-called edge-of-stability phase. During this phase, the Hessian of the loss function remains approximately constant close to the stability threshold $\frac{2}{\eta}$. Meanwhile, the loss is still decreasing in the long run, but no longer monotonically. Cohen et al. (2020) also studied whether stochastic gradient descent exhibits the same behavior. In Appendix G, they observe that for stochastic gradient descent progressive sharpening stops before the threshold $\|\operatorname{Hess}\mathcal{L}(X_{n}^{\operatorname{GD}})\|\approx\frac{2}{\eta}$, is reached, suggesting that SGD does not enter the edge of stability. In Appendix H, however, they observe that the expected change of the loss does behave similarly to the non-monotonous decrease observed for gradient descent at the edge of stability. Based on this, they conjecture (cf. Section 6 in Cohen et al., 2020) that SGD does enter an edge-of-stability regime, except that $\|\operatorname{Hess}\mathcal{L}(x)\|\leq\frac{2}{\eta}$ is no longer the correct notion of stability for gradient descent (cf. also the discussions in Wu et al., 2018 and Andreyev and Beneventano, 2025).
Instead of directly tracking the sharpness of the loss, $\|\operatorname{Hess}\mathcal{L}(x)\|$, one can equivalently study the evolution of $\mu(x)$ along the trajectory. It should be mentioned that in Section 2.3 we have only defined $\mu(x)$ for the parameters $x$ that lie in the manifold of interpolation solutions $\mathcal{M}$. However, the equivalent definition (2.18) of $\mu$ given in Section 2.5, is well defined for any $x\in\mathbb{R}^{D}$. Thus, it is possible to track $\mu(X_{n}^{\operatorname{GD}})$ during training. Since $\mu$ is related to the Hessian of the loss by $\mu(x)=\log\big{|}1-\eta\|\operatorname{Hess}\mathcal{L}(x)\|\big{|}$, it contains the same information as $\|\operatorname{Hess}\mathcal{L}(x)\|$. In the experiments of Cohen et al. (2020), we would observe that $\mu(X_{0}^{\operatorname{GD}})$ is negative and that $\mu(X_{n}^{\operatorname{GD}})$ progressively increases until it enters the edge of stability $\mu(X_{n}^{\operatorname{GD}})=0$ from which point on $\mu(X_{n}^{\operatorname{GD}})$ remains close to zero. Based on our main results, we suggest that from a dynamical point of view, tracking $\lambda(X_{n}^{\operatorname{SGD}})$ during training with stochastic gradient descent is the appropriate analogue to the study of Cohen et al. (2020). We leave it as an open problem to the community to determine the behavior of $\lambda$ by the means of empirical experiments. An interesting challenge in conducting such a study lies in the numerical computation of the Lyapunov exponent for high-dimensional random-matrix products. Numerical schemes for the computation of Lyapunov exponents can be found, for example, in Eckmann and Ruelle (1985) and Sandri (1996).
@@ 2.7 Possible Extensions and Outlook
While our main results are general in the sense that they require only very weak assumptions on the network function $\mathfrak{F}$, we want to point out that we only consider scalar regression problems and training with gradient descent or stochastic gradient descent with mini-batch size 1. As mentioned above, we have made these restrictions to simplify the already extensive proofs as much as possible. The main results can be extended, with minor modifications, to a more general setting.
In practice, SGD is usually implemented with so-called mini-batches to allow for parallel computations. For SGD with mini-batches of size $1\leq B\leq N$, the iterative step (2.8) is replaced by
@@
where $G_{x^{*},[\Xi]}$ denotes the Gram matrix $G_{x^{*}}$ of the neural tangent kernel with all rows whose index is not in $\Xi$ set to 0. In order to derive a version of Theorem B it is necessary to adapt the notion of regular global minima (cf. Definition 2.2) to ensure that the analog of Lemma 3.9 in the proof below still remains valid. We leave this as a problem for future work. The rest of the proof of Theorem B can be extended to mini-batch SGD with only minor changes. Such an extension of Theorem B could help to better understand the impact of the learning rate and the mini-batch size on generalization properties (see e.g. Hoffer et al., 2017; Goyal et al., 2017; Keskar et al., 2017 for numerical studies).
@@
Furthermore, it would be interesting to extend our analysis to other optimizers such as SGD with momentum (cf. Rumelhart et al., 1986) or Adam (cf. Kingma and Ba, 2014). The derivations in Section 2.3 can be extended to these algorithms by linearizing the iteration steps around fixed points where all moment terms are zero. In the case of Adam, the parameter $\epsilon$ appearing in the numerator of the final update step is usually chosen extremely small (Kingma and Ba, 2014 suggest $\epsilon=10^{-8}$). Thus, the linearization only approximates the dynamics of the actual algorithm in a vanishingly small neighborhood, and it is questionable whether it is still meaningful for the dynamics of the algorithm.
@@
Our analysis can be extended to multidimensional regression problems with ground truth function $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}^{\tilde{d}}$ and network function $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}^{\tilde{d}}$. In fact, one can interpret such a multidimensional regression problem as a scalar regression problem with a ground truth function $\hat{f}:\mathbb{R}^{d}\times[\tilde{d}]\to\mathbb{R}$ and a network function $\mathfrak{F}:\mathbb{R}^{D}\times(\mathbb{R}^{d}\times[\tilde{d}])\to\mathbb{R}$. This has the effect of splitting one training example $(y,z)\in\mathbb{R}^{d}\times\mathbb{R}^{\tilde{d}}$ into $\tilde{d}$ training examples $((y,1),z_{1}),\dots,((y,\tilde{d}),z_{\tilde{d}})\in(\mathbb{R}^{d}\times[\tilde{d}])\times\mathbb{R}$. An SGD step with the training example $(y,z)$ now corresponds to an SGD step with the mini-batch $((y,1),z_{1}),\dots,((y,\tilde{d}),z_{\tilde{d}})$. Similarly to mini-batch SGD, the main challenge in extending Theorem B to multidimensional regression problems is to adapt the notion of a regular global minimum (cf. Definition 2.2) in such a way that Lemma 3.9 still remains valid for multidimensional regression.
@@
Unfortunately, our results cannot be easily extended to classification problems. For a classification task with $k$ classes, it is common to encode the training data as pairs $(y,z)\in\mathbb{R}^{d}\times\mathbb{R}^{k}$, where $z=e_{i}$ is the $i$-th unit vector, where $i\in[k]$ is the class $y$ belongs to. Then one considers a network function $\mathfrak{F}:\mathbb{R}^{D}\times\mathbb{R}^{d}\to\mathbb{R}^{k}$, which is then postcomposed with the softmax function $\operatorname{softmax}_{\beta}:\mathbb{R}^{k}\to[0,1]^{k}$ given by
@@
where $\beta>0$ is some positive real number commonly referred to as inverse temperature. For training, the most common loss function is the cross-entropy loss given by
@@
To have $\ell(\bar{z},z)=0$, one necessarily needs $\bar{z}=z$. However, the outputs of the softmax function are in $(0,1)^{k}$ for finite input values. Thus, it is not possible to reach training error zero and $\mathcal{M}=\emptyset$.
Finally, it should be mentioned that our analysis assumes that the model is trained to convergence. In practice, it has been observed that stopping the algorithm early can improve generalization. While the global minima with $\lambda(x^{*})>0$ are asymptotically unstable, the finite time Lyapunov exponents
@@
can, even for large $n$, still be negative with small, but positive probability. This effect is captured by a large deviation principle (cf. Arnold and Kliemann, 1987), the rate function of which depends on the moment Lyapunov exponents (see Definition A.2).